ment CALCULATOR MESSAGE MY INSTRUCTOR FULL SCREEN PRINTER VERSION BACK NEXT Do It! Review 11-3a Cheyenne Corp. has 2,400 shares of 9%, $101 par value preferred stock outstanding at December 31, 2017. At December 31, 2017, the company declareda $119,000 cash dividend Determine the dividend paid to preferred stockholders and common stockholders under each of the following scenarios 1. The preferred stock is noncumulative, and the company has not missed any dividends in previous years The dividend paid to preferred stockholders s The dividend paid to common stockholders 2. The preferred stock is noncumulative, and the company did not pay a dividend in each of the two previous years. The dividend paid to preferred stockholders s The dividend paid to common stockholders 3. The preferred stock is cumulative, and the company did not pay a dividend in each of the two previous years. The dividend paid to preferred stockholders The dividend paid to common stockholders s would like to Show Wor INK TO TEXT
Answer #1

1) Dividend distribution :

The dividend paid to preferred stockholders = 2400*101*9% = 21816

The dividend paid to Common stockholders = 119000-21816 = 97184

2) Dividend distribution :

The dividend paid to preferred stockholders = 2400*101*9% = 21816

The dividend paid to Common stockholders = 119000-21816 = 97184

3) Dividend distribution :

The dividend paid to preferred stockholders = 2400*101*9%*3 = 65448

The dividend paid to Common stockholders = 119000-65448 = 53552